Determine the mathematical relationship that exist among T, L and W? If a sample of DNA twists around the duplex axis six times, and it contains one negative and one positive supercoil which was relieved by a topoisomerase, what is the Linking number?
Answer #1

Lk  = T + W

Lk - linking number

T - no. of complete turns one strand makes

W - no. of times dulex axis cross

Relaxed DNA negatively supercoiled positively supercoiled

Topoismorase uses energy to change the linking number from the base number. Increase in linkage called positive recoiling and decrease in linkage called negative recoiling.

In relaxed DNA , W = 0, so L = T= bp/ 10

In positively coiled, W = 5 , T = 6, so, L = T+ W = +11

In negatively coiled, W = - 5, T = 6 So, L = T +W = +1